Using either substitution or elimination of the following system solve 6x+(3÷2)y=-12 and -4x-y=4.
Answer by Alan3354(69443)

3y/2 =-12
-4x-y =4
----------
3y/2 =-12
3y = -24
y = -8
============
-4x-y =4
-4x + 8 = 4
-4x = -4 x = 1